OpenStudy (anonymous):

all polynomial a+a1x+a2x^2+a3x^3 for which a+a1+a2+a3=0 is this a subspace of p3

OpenStudy (aroub):

p3?

OpenStudy (anonymous):

p3 shows polynomial3

OpenStudy (anonymous):

As in the previous problem you need to see if it satisfies all of the vector space axioms. A subspace must have a zero vector and must be closed under addition and scalar multiplication.

OpenStudy (anonymous):

In order to show that the set S of cubic polynomials, the sum of whose coefficients os 0, is a subspace, we must show that S is closed under addition. If we find even one counterexample, i.e., two polynomials in S whose sum is not in S, that is sufficient to prove S is not a subspace of \[p _{3}.\]

OpenStudy (anonymous):

1) is the 0 polynomial in the space? well, 0+0x+0x^2+0x^3 is the 0 polynomial, 0+0+0+0 = 0, so yes, it is in the space :)

OpenStudy (anonymous):

Just looking at the coefficients, you can see that its closed under addition. a_1 + a_2 + ... = 0 b_1 + b_2 + ... = 0 a_1 + a_2 + .. + b_1 + b_2 + .. = 0

OpenStudy (anonymous):

2) if :\[a+a_1x+a_2x^{2}+a_3x^{3}, b+b_1x+b_2x^{2}+b_3x^{3} \] are 2 polynomials in the space, is their sum still in the space? adding them gives: \[(a+b)+(a_1+b_1)x+(a_2+b_2)x^{2}+(a_3+b_3)x^{3}\] where \[(a+b)+(a_1+b_1)+(a_2+b_2) + (a_3+b_3) = (a+a_1+a_2+a_3)+(b+b_1+b_2+b_3) = 0+0 = 0\] so it is still in the space, and is closed under addition.

OpenStudy (anonymous):

oops, got cut off >.< last lines are 0 + 0 = 0

OpenStudy (anonymous):

Also obviously closed under scalar multiplication.

OpenStudy (anonymous):

3) if: \[a+a_1x+a_2x^2+a_3x^{3} \] is in the space, and i multiply it by a scalar, is it still in the space? multiplying by c we get: \[ca+ca_1x+ca_2x^2+ca_3x^{3}\] where: \[ca+ca_1+ca_2+ca_3 = c(a+a_1+a_2+a_3) = c(0) = 0\] so its closed under scalar multiplication.